What is the twenty-ninth term of arithmetic sequence with a1=13 and d=-5/2?
A)-62
B)83
C) 88
D) -57

Answer by richwmiller(17219)

an = a + (n - 1)d
a29 = 13 + (28)-5/2
a29 = 13 + (14)-5
a29 = 13 + -70
a29=-57
D) -57